The Stewart Company has $2,286,500 in current assets and $937,465 in current liabilities. Its initial inventory level is $640,220, and it will raise funds as additional notes payable and use them to increase inventory. How much can its short-term debt (notes payable) increase without pushing its current ratio below 2.0? Round your answer to the nearest cent.

$  
Answer #1
The current ratio shouldn't go below 2, it means that maximum reduction in current ratio up to 2 is acceptable, so the desired current ratio is 2
Let us assume X be the amount of short term debt (notes payable) which will be used to increase the inventory.
Desired current ratio = ( Current assets + X ) / ( Current liabilities + X )
2 = ( 2286500 + X ) / ( 937465 + X )
2 * ( 937465 + X ) = 2286500 + X
1874930 + 2X = 2286500 + X
2X - X = 2286500 - 1874930
X = 2286500 - 1874930 411570.00